# The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives

### v3

@@ 5.2 LLMs for Content Moderation
Table 5 shows the Accuracy and False Positive Rate (FPR) for content classification for each model. Our aim is to minimize the FPR because our priority is to avoid mistakenly labeling inappropriate stories as appropriate, as this could result in children being exposed to unsuitable content. Ensuring high accuracy with a low FPR is essential for maintaining a safe and child-friendly storytelling environment. GPT-4o has the lowest FPR of 9%. The model was tested on human-written stories, where the risk of inappropriate content is higher. However, when generating content, LLMs have built-in safeguards (Wang et al., 2024) that significantly reduce the probability of producing such inappropriate material, making the likelihood of harmful content in generated stories very low.
Reference Test Model Versus Rater 1 Rater 2
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
Female XTTSv2†‡ StyleTTS 2 48 24 28 62 18 20
Male XTTSv2†‡ StyleTTS 2 50 28 22 46 32 22
Table 4: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
Models Accuracy (%) FPR (%)
GPT-4o-mini 73 26
GPT-4o 90 9
Llama-3.1-8b 59 39
Llama-3.1-70b 67 32
Gemma-2-9b 65 34
Gemma-2-27b 57 42
Table 5: Accuracy and False Positive Rate (FPR) for content classification.
@@ 5.3 Text-to-Speech
We calculated the weighted Cohen’s Kappa, as shown in Table 6, to assess the agreement between total scores by the two raters. The results indicate good inter-rater agreement (Kraemer, 2015), with almost all Kappa scores greater than 0.6.
Female Reference Male Reference
XTTSv2 StyleTTS 2 XTTSv2 StyleTTS 2
$\kappa$ 0.601 0.579 0.644 0.701
Table 6: $\kappa$ between the two human rater for TTS models.
Table 4 shows the comparative performance of two TTS models, XTTSv2 and StyleTTS 2, evaluated on both female and male reference audios by two annotators.
For female reference audio, based on the scoring done by Rater 1, XTTSv2 has a win-rate of 48%, while StyleTTS 2 had a win-rate of 28% and the tie-rate is 24%. This indicates that XTTSv2 was generally favored by Annotator 1, though a significant portion of the comparisons resulted in ties. Annotator 2, however, showed a much stronger preference for XTTSv2, with a win-rate of 62%, and only an 18% tie-rate, suggesting greater confidence in the performance of XTTSv2 for female voice. Figure 3 shows that StyleTTS 2 outperforms XTTSv2 in only Fluency and Pronunciation for female voice.
Figure 3: Comparison of XTTSv2 and StyleTTS 2 (female voice) for each metric for both raters.
For male reference audio, Annotator 1’s scores gave XTTSv2 a win-rate of 50%, demonstrating an edge over StyleTTS 2, which had a win-rate of 22%. Based on Annotator 2’s evaluation of male reference audio, XTTSv2 won 46% of the time and tying 32% of the time. This indicates that, while XTTSv2 performed better, StyleTTS 2 was considered slightly more competitive with male voice. Figure 4 shows that StyleTTS 2 outperforms XTTSv2 in Fluency and Pronunciation according to both raters for male voice. According to Rater 2 StyleTTS 2 is slightly better in Clarity and Pauses category.
Figure 4: Comparison of XTTSv2 and StyleTTS 2 (male voice) for each metric for both raters.
Overall results showed that XTTSv2 outperformed StyleTTS 2, where it achieved higher win-rates and lower loss-rates across both annotators and reference audios. Both models were given a voice cloning score of 1 out of 2 by both annotators. Based on the results of the evaluation we use XTTSv2 for our system.
@@ 5.4 Text-to-Video
We calculated the weighted Cohen’s Kappa, as shown in Table 7, to assess the agreement between total scores by the two raters. The results indicate good inter-rater agreement, with all Kappa scores greater than 0.6.
Cartoon Anime Animated
$\kappa$ 0.616 0.626 0.635
Table 7: $\kappa$ between the two human rater for TTV model.
Test Style Versus Rater 1 (%) Rater 2 (%)
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
Cartoon Anime†‡ 20 26 54 26 16 58
Cartoon Animated 50 12 38 26 30 44
Anime†‡ Animated 54 30 16 34 34 32
Table 8: Win-rate, tie-rate, and loss-rate for style 1 against style 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
The results in Table 8 present a detailed comparison of the performance between three video styles — Cartoon, Anime, and Animated — based on evaluations by two raters. The table quantifies how often one style outperformed, tied, or lost to another in a pairwise comparison across all evaluated videos.
When comparing Cartoon and Anime, Rater 1 observed that Cartoon won only 20%, tied 22%, and lost 56% of the comparisons to Anime. Similarly, Rater 2 reported that Cartoon won in 26% of the instances, tied in 16%, and lost in 58%. This shows that Anime consistently outperformed Cartoon according to both raters, with only a slight variation in the win and tie rates between the two.
In the comparison between Cartoon and Animated, Rater 1 found that Cartoon performed significantly better, winning 50% of the time, tying in 12%, and losing in 38% of cases. Rater 2’s evaluation, however, showed that Cartoon won 26% of the comparisons, tied in 30%, and lost in 44% of them. This indicates that while Rater 1 favored Cartoon over Animated, Rater 2 observed Animated style slightly outperforming Cartoon style overall.
Finally, in the comparison between Anime and Animated, both raters favored Anime. Rater 1 found that Anime won 54% of the time, tied in 30%, and lost only 16% of the comparisons to Animated. Rater 2 reported Anime winning 34%, tying in 34%, and losing in 32%. Despite some differences in the magnitude of win, tie, and lose rates, both raters agreed that using Anime style generated better video quality.
Figure 5 shows that all three styles have similar performance in Overall Alignment, Fine-Grained Alignment, and Child Friendliness. However, in both evaluations, Anime outperforms other styles in Temporal Quality and Naturalness assessment by a big margin. Based on the results of the evaluation we use Anime style animation for our system.
@@ 5.5 Final Design
Based on extensive evaluation of each component of our system, we selected the following models: for the Writer, we use Llama-3.1-8b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogXVideo-5b is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
@@ 7 Future Work
In future work, several enhancements could be made to improve the system’s capabilities and user experience. One key area is the integration of image inputs, where drawings provided by children could be used as prompts for story generation. Additionally, incorporating different voices for distinct character dialogues could greatly enhance the immersion of the generated audio-visual stories. Ensuring consistency in the characters and environment across all generated videos is another important direction. Furthermore, adding AudioGen (Copet et al., 2023) as the Foley Artist agent to incorporate sound effects based on the video content would bring the stories to life. These advancements have the potential to significantly enhance the overall interactivity and immersion of the system.
Acknowledgments
We are grateful to our Research Assistants: Sehar Salman Haider and Syeda Duaa Batool Naqvi who were involved in human evaluation of the data.
Appendix A System Screen
@@
Appendix B Score Breakdown
@@
Appendix C System Prompts
@@
Figure 6: Frontend design for the system.
@@
Fluency and Pronunciation 0: Pronunciation is correct and speech is fluent.
@@
Voice Cloning 0: The cloned voice is almost identical to the original, with a consistent tone, pitch, and speech pattern, making it indistinguishable from the intended voice.
@@
Temporal Quality 0: Transitions between frames are smooth and coherent, with no jarring jumps or stutters.
@@
2: Major misalignments, such as wrong colors, incorrect motion directions, or speeds that don’t match the description.

### v4

@@ 5.2 LLMs for Content Moderation
Models Gutenberg Synthetic
GPT-4o-mini 26 25
GPT-4o 9 0
Llama-3.1-8b 39 37.5
Llama-3.1-70b 32 37.5
Gemma-2-9b 34 0
Gemma-2-27b 42 12.5
Table 3: False Positive Rate (FPR) for content classification.
Table 3 shows the False Positive Rate (FPR) for content classification for each model. Our aim is to minimize the FPR because our priority is to avoid mistakenly labeling inappropriate stories as appropriate, as this could result in children being exposed to unsuitable content. Ensuring a low FPR is essential for maintaining a safe and child-friendly storytelling environment. GPT-4o achieves the lowest FPR, with 9% FPR for the Project Gutenberg dataset and 0% FPR for the LLM-generated story dataset.
The Project Gutenberg dataset contains human-written stories, where the risk of inappropriate content is higher. However, when generating content, LLMs have built-in safeguards (Wang et al., 2024) that significantly reduce the probability of producing inappropriate material, making the likelihood of harmful content in generated stories very low.
@@ 5.3 Text-to-Speech
Table 4 shows how the two text-to-speech models—XTTSv2 and StyleTTS 2—fared under female and male reference voices. When using the female reference audio, XTTSv2 achieves a 37.83% win-rate against StyleTTS 2 but also has a slightly higher loss-rate of 39.00%, reflecting a very close quality. By contrast, with the male reference audio, XTTSv2 performs somewhat better overall (39.67% win vs. 34.67% loss). Tie-rates remain in the 23–26% range for both scenarios, indicating that both models produce similarly acceptable outputs in a substantial fraction of cases.
Although these comparisons do not reveal a single dominant TTS system across all conditions, the trends suggest that XTTSv2 may slightly outperform StyleTTS 2 in handling the male voice, whereas StyleTTS 2 holds a marginal edge—or at least near parity—when cloning the female voice.
Rates (%)
Test Model Versus Model Win Tie Loss
XTTSv2 StyleTTS 2 37.83 23.17 39.00
XTTSv2 StyleTTS 2 39.67 25.67 34.67
Table 4: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. Row 1 is for female reference audio and row 2 is for male reference audio.
Figure 3 shows a category-wise breakdown of each model’s average TTS scores using a female reference voice. When evaluated with a female reference voice, StyleTTS 2 achieves higher scores in Clarity and Pauses (1.68 vs. 1.60) as well as Fluency and Pronunciation (1.88 vs. 1.75), suggesting it provides more precise articulation and smoother pacing overall. Although StyleTTS 2 performs better in Voice Cloning (1.00 vs. 0.17), it score is still low. By contrast, XTTSv2 scores notably higher in Information and Emotion Preservation (1.61 vs. 1.23), implying it better conveys the emotional cues. In terms of Intonation and Naturalness, the two models are relatively close (1.53 vs. 1.40), though XTTSv2’s slightly higher score points to marginally more human-like pitch variation. Overall, these results suggest that StyleTTS 2 may be the stronger choice when precise pronunciation and voice replication are prioritized, whereas XTTSv2 appears better suited for scenarios that demand heightened emotional expressiveness.
Figure 3: Comparison of XTTSv2 and StyleTTS 2 (female voice) for each metric.
Figure 4 provides a category-wise breakdown of each model’s average scores for text-to-speech tasks using a male reference voice. XTTSv2 and StyleTTS 2 both demonstrate relatively high average scores in Clarity and Pauses (1.63 vs. 1.67), reflecting minimal listening effort and well-paced output. However, XTTSv2 scores noticeably higher in Information and Emotion Preservation (1.70 vs. 1.32) and Intonation and Naturalness (1.56 vs. 1.46), suggesting it captures more nuanced expressive cues and conveys pitch and rhythm changes in a more human-like manner. In contrast, StyleTTS 2 excels in Fluency and Pronunciation (1.85 vs. 1.77) and Voice Cloning (0.67 vs. 0.50). Overall, the choice between these two TTS systems lies on whether rich emotional conveyance (XTTSv2) or speaker-identity matching (StyleTTS 2) takes priority for a particular application.
Figure 4: Comparison of XTTSv2 and StyleTTS 2 (male voice) for each metric.
@@ 5.4 Text-to-Video
Table 5 summarizes the relative performance of three text-to-video animation styles—Cartoon, Anime, and Animated—based on head-to-head comparisons. Cartoon lags behind Anime (28.67% win vs. 48.33% loss) and also against Animated (28.33% win vs. 49.50% loss), suggesting that its visuals are generally of lower quality. On the other hand, Animated leads when paired with Anime (34.33% vs. 36.00% loss), although its margin is not overwhelmingly large. Tie-rates hover around 20–30%, reflecting that all three styles can produce some visually appealing results under certain prompts. Overall, Animated style appears slightly stronger than both Cartoon and Anime in crafting engaging visuals, whereas Cartoon tends to be the weakest performer in pairwise comparisons.
Rates (%)
Test Model Versus Model Win Tie Loss
Cartoon Anime 28.67 23.00 48.33
Cartoon Animated 28.33 22.17 49.50
Anime Animated 34.33 29.67 36.00
Table 5: Win-rate, tie-rate, and loss-rate for style 1 against style 2 based on scoring by both raters.
Figure 5 shows a category-wise breakdown of each animation style’s average scores for text-to-video tasks. When evaluating different animation styles based on five criteria, the Animated style yields the strongest Overall Alignment score (1.60) and highest Child-Friendliness (1.80). By contrast, Anime achieves the top marks for both Naturalness Assessment (1.36) and Temporal Quality (1.44), suggesting that it often offers smoother transitions and fewer visual artifacts. Although Cartoon generally scores lower in Naturalness Assessment (1.18) and Temporal Quality (1.33), it remains reasonably competitive in Overall Alignment (1.49) and offers decent Child-Friendliness (1.70). Overall, Animated stands out for aligning the generated visuals with text prompts, while Anime excels in natural-feeling motion and coherent scene transitions; Cartoon, meanwhile, appears to lag somewhat behind the other two styles, though it retains a solid score for child-centric content.
@@ 5.5 Final Design
Based on extensive evaluation of each component of our system, we selected the following models: for the Writer, we use Gemma-2-9b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogXVideo-5b, with Animated style, is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
@@ 7 Future Work
In future work, several enhancements could be made to improve the system. One key area is the integration of image inputs, where drawings provided by children could be used as prompts for story generation. Additionally, incorporating different voices for distinct character dialogues could greatly enhance the immersion of the generated audio-visual stories. Furthermore, adding AudioGen (Copet et al., 2023) as the Foley Artist agent to incorporate sound effects based on the video content would bring the stories to life.
8 Limitations
The system has several limitations that need to be addressed to enhance its performance and effectiveness. One major limitation is the need for a more comprehensive dataset to robustly test the child-appropriateness of the generated content. Additionally, the generated animations, while not inappropriate, often suffer from visual distortions and lack the smoothness and coherence required for engaging and child-friendly storytelling. These visual inconsistencies can detract from the immersive experience and may not fully align with the intended narrative quality. Another significant limitation is the high generation time for animations, with each six-second video requiring approximately three minutes to render.
9 Ethical Impact
The ethical impact of this system centers on ensuring child safety and fostering positive developmental outcomes. By integrating robust content moderation and focusing on child-appropriate narratives, the system aims to provide a safe storytelling environment. However, ethical considerations include the potential for cultural biases in generated content, requiring careful dataset curation and evaluation to ensure inclusivity and fairness. Additionally, the use of generative AI must prioritize transparency, explicitly informing users about AI-generated content to maintain trust. Addressing these ethical concerns is critical to creating a responsible and impactful educational tool for children.
Appendix A Implementation Details
A.1 Models
Llama-3.1 models are available under llama3.1 license and Gemma-2 models are under gemma license. GPT-4 is available under proprietary licence. For TTS models, XTTSv2 is under coqui-public-model-license and StyleTTS 2 is using MIT license. The TTV model, CogVideoX is using it’s own custom license. Finally, MusicGen model uses CC-BY-NC-4.0. All models used in this paper comply with their respective license.
A.2 Dataset
Our study adheres to the Project Gutenberg License, which allows the use of its texts for non-commercial purposes with proper attribution, ensuring compliance with their terms and conditions.
A.3 Model Size and Budget
We use Llama-3.1 in two sizes: 8 billion and 70 billion parameters, and Gemma-2 in two sizes: 9 billion and 27 billion parameters. A single Nvidia A100 80GB GPU was used to deploy the full system.
A.4 Human Annotators
There are six human annotators in this study, all of them are undergraduate Computer Science students. All annotators are native speakers of Urdu from Pakistan with English as the medium of education.
Appendix B System Screen
@@
Appendix C Score Breakdown
@@
Appendix D System Prompts
@@
Figure 6: The system’s front-end design.
@@
Fluency and Pronunciation 0: Pronunciation is inaccurate and speech is not comprehensible.
@@
Voice Cloning 0: The cloned voice is not recognizable, with noticeable differences in tone, pitch, or speech pattern, and can not be identified as the intended voice.
@@
Temporal Quality 0: Transitions between frames are abrupt and incoherent, with noticeable stutters and jarring jumps.
@@
2: All attributes align with the prompt, with no discrepancies.